# AutoCrawler : A Progressive Understanding Web Agent for Web Crawler Generation

### v1

@@ 3 AutoCrawler
In this section, we describe our framework AutoCrawler for generating a crawler to extract specific information from semi-structured HTML. Our approach is divided into two phases: first, we adopt a progressive generation framework that utilizes the hierarchical structure of web pages; second, we employ a synthesis framework based on results from multiple web pages. The overall framework is presented in Figure 2.
@@ 18       $k\leftarrow k+1$;
3.1 Modeling
Unlike the wrapper method that generates an XPath, we model the crawler generation task as an action sequence generation task. In specific, we generate an action sequence $\mathcal{A}_{seq}$ that consists of a sequence of XPath [^4] https://en.wikipedia.org/wiki/XPath expression from a set of seed webpages (i.e., a small portion of webpages in the test case for generating the sequence).
Models Method Executable Evaluation IE Evaluation
Closed-source LLMs
GPT-3.5-Turbo COT 36.75 8.83 6.71 43.46 0.71 3.53 89.45 50.43 47.99
Reflexion 46.29 11.66 2.83 37.10 0.71 1.41 94.67 55.85 55.10
AutoCrawler 54.84 11.83 8.96 19.35 1.08 3.94 85.85 73.34 69.20
[…]
Reflexion 33.12 6.56 4.06 52.50 0.63 3.12 87.45 42.75 40.88
AutoCrawler 42.81 11.87 4.69 34.38 1.25 5.00 85.70 57.54 54.91
[…]
Reflexion 67.50 13.75 4.37 10.94 0.94 2.50 93.28 82.76 82.40
AutoCrawler 71.56 14.06 5.31 4.06 0.63 4.37 92.49 89.13 88.69
Open-source LLMs
Mistral 7B COT 3.44 0.31 0.63 95.31 0.00 0.63 94.23 4.55 4.24
Reflexion 2.19 0.00 0.31 97.19 0.00 0.31 95.60 2.78 2.49
AutoCrawler 2.87 0.00 0.00 96.77 0.36 0.00 98.57 3.23 2.87
[…]
Reflexion 18.08 4.80 2.95 73.06 0.00 1.11 78.96 23.26 22.44
AutoCrawler 23.99 8.12 1.48 64.94 0.00 1.48 78.59 28.70 28.41
[…]
Reflexion 36.25 6.88 3.12 51.25 0.00 2.50 89.35 44.57 43.60
AutoCrawler 46.88 10.62 7.19 30.31 0.63 4.37 87.32 62.71 59.75
[…]
Reflexion 37.19 11.25 4.06 44.69 1.25 1.56 86.41 48.28 47.08
AutoCrawler 38.75 11.25 5.31 39.69 0.63 4.37 84.91 52.11 49.68
Table 1: The executable evaluation and IE evaluation of LLMs with three frameworks in Swde dataset. We examine 7 LLMs, including 3 closed-source LLMs and 4 open-source LLMs.
3.2 Progressive Generation
Dealing with the lengthy content and hierarchical structure of webpages, generating a complete and executable crawler in one turn is difficult. However, the HTML content is organized in a DOM tree structure, which makes it possible to prune irrelevant page components and hence, limit the length and height of the DOM tree to improve the performance of LLM generation.
3.3 Synthesis
Specifically, we randomly select $n_{s}$ webpages from the test case as seed webpages. Then, we generate an action sequence for each of them. Subsequently, we execute multiple different action sequences to extract information from the seed webpages, respectively. We collect all action sequence and their corresponding results and then choose one action sequence that can extract all the target information in the webpages as the final action sequence.
4 Experiment
Intending to put AutoCrawler to practical use, we investigate the following research questions: 1) How is the overall performance of AutoCrawler in comparison to the current state-of-the-art crawler generation task? 2) How does our progressive understanding generation framework improve performance? What is its relationship to the size of LLM? 3) Can we rely entirely on LLMs for web automation? 4) In which scenarios do current frameworks still not perform well?
4.1 Experimental Settings & Evaluation Metrics
We conduct our experiment on various LLMs including closed-source LLMs: GPT-3.5-Turbo OpenAI (2022), Gemini ProTeam et al. (2023) and GPT4 OpenAI (2023) as well as open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al. (2023), CodeLlama-34B-Instruct Rozière et al. (2024), Mixtral 8$\times$7B-Instruct-v0.1 Jiang et al. (2024) and Deepseek-Coder-33B-Instruct Guo et al. (2024). Furthermore, we apply different LLM-prompt-based web agents as our baselines, including ZS_COT Wei et al. (2023) and Reflexion Shinn et al. (2023) and AutoCrawler to them. Due to the limited-length context of LLMs, all experiments are conducted under zero-shot settings. The full prompts can be found in Appendix D.2.
We test them on three datasets: Swde Hao et al. (2011), Extend Swde Lockard et al. (2019) and Ds1 Omari et al. (2017). The experimental result of the last two can be found in Appendix A.1 and  A.2. We set the size of seed webpages $n_{s}=3$ and max retry times $d_{max}=5$.
In addition to the execution evaluation metrics described in Section 2.3, we also employ traditional evaluation metrics to more comprehensively assess the quality of different action sequences. Specifically, we adopt precision (P.), recall (R.), and macro-f1 (F1), which are calculated as the mean of the corresponding metrics for each case.
4.2 Main Results on Swde
Results in Table 3.1 show that: 1) With AutoCrawler generating action sequence, LLMs can achieve better performance. Compared to the COT and Reflexion baseline, our method performs a higher ratio of correct and lower ratio of un-executable. Also, it should be noted that Mixtral 8$\times$7B + AutoCrawler can outperform ChatGPT + Reflexion, indicating the superiority of AutoCrawler in the generation of executable action sequences in the crawler generation task. 2) Models with small parameter sizes have significant difficulties in understanding and writing executable paths, so they can be considered challenging to apply in this task. On the contrary, large-scale models demonstrate a more stable ability in instruction alignment, web structure comprehension, and reflection on execution results; 3) Traditional IE evaluation metrics cannot well describe the success rate of our task. Especially for the precision metric, it fails to reveal the performance gap among different methods with different models. This is because the extraction metrics only evaluate the results that have been extracted, ignoring that unexecutable or empty extractions also greatly damage the executability.
4.3 Generate with Golden Label
Models Method Executable Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else
Closed-source LLMs
GPT-3.5-Turbo COT 41.70 12.92 7.38 35.42 0.74 1.85
Reflexion 47.23 16.24 2.21 33.21 0.37 0.74
AutoCrawler 56.89 19.43 5.65 13.43 0.71 3.89
[…]
Reflexion 35.31 9.38 6.88 43.75 1.56 3.12
AutoCrawler 45.31 13.44 6.25 30.31 1.25 3.44
[…]
Reflexion 71.25 7.19 4.69 14.37 0.94 1.56
AutoCrawler 75.31 10.94 4.37 4.06 0.63 4.69
Open-source LLMs
Mistral 7B COT 2.19 0.00 0.31 97.19 0.00 0.31
Reflexion 2.19 0.00 0.00 97.50 0.31 0.00
AutoCrawler 2.19 0.00 0.00 97.19 0.31 0.31
[…]
Reflexion 22.21 4.93 3.94 66.95 0.49 1.48
AutoCrawler 26.20 12.55 5.54 53.51 0.00 2.21
[…]
Reflexion 34.69 8.13 5.31 49.06 0.63 2.19
AutoCrawler 45.62 11.56 5.94 32.50 1.25 3.12
[…]
Reflexion 38.75 11.87 2.81 42.19 0.63 3.75
AutoCrawler 38.44 20.94 4.06 31.56 0.94 6.56
Table 2: The executable and IE evaluation with 7 LLMs on Swde dataset with golden label.
To better illustrate the effectiveness of our framework in generating executable action sequences, we compare the performance of the COT, Reflexion, and our framework, while giving the golden label of the instruction. By offering the same extraction targets, we can effectively detect the actual effects of different frameworks on generating action sequences.
Table 4.3 shows experimental results, from which we can have the following observations: 1) Our proposed progressive understanding framework still effectively enhances the model’s performance under this setting; 2) LLMs still suffer in accurately understanding web page contents with semi-structured markup languages, which illustrate the performance gap between Table 3.1 and Table 4.3; 3) Compared to closed-source LLMs, even provided with golden labels, Open-source LLMs are unable to achieve sustained performance improvement. This phenomenon demonstrates that the bottleneck for these models lies not in understanding the webpage content but in understanding the webpage’s hierarchical structure itself.

### v2

@@ 4 AutoScraper
In this section, we describe our framework AutoScraper for generating a scraper to extract specific information from semi-structured HTML. Our approach is divided into two phases: first, we adopt a progressive generation module that utilizes the hierarchical structure of web pages; second, we employ a synthesis module based on results from multiple web pages. The overall framework is presented in Figure 2.
4.1 Modeling
Unlike the wrapper method that generates an XPath, we model the scraper generation task as an action sequence generation task. In specific, we generate an action sequence $\mathcal{A}_{seq}$ that consists of a sequence of XPath [^5] https://en.wikipedia.org/wiki/XPath expression from a set of seed webpages (i.e., a small portion of webpages in the test case for generating the sequence).
Dataset Swde Extended Swde Ds1
Models Method Executable Evaluation IE Evaluation Exec Eval IE Eval Exec Eval IE Eval
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1 Correct Unex. F1 Correct Unex. F1
Closed-source LLMs
GPT-3.5-Turbo COT 36.75 8.83 6.71 43.46 0.71 3.53 89.45 50.43 47.99 35.19 55.40 41.28 32.65 53.06 41.16
Reflexion 46.29 11.66 2.83 37.10 0.71 1.41 94.67 55.85 55.10 43.90 49.13 48.66 36.73 51.02 43.75
AutoScraper 54.84 11.83 8.96 19.35 1.08 3.94 85.85 73.34 69.20 46.34 34.84 57.74 48.98 44.90 52.38
[…]
Reflexion 33.12 6.56 4.06 52.50 0.63 3.12 87.45 42.75 40.88 34.15 51.57 41.66 20.25 65.82 27.66
AutoScraper 42.81 11.87 4.69 34.38 1.25 5.00 85.70 57.54 54.91 35.89 42.86 47.80 43.04 34.18 56.92
[…]
Reflexion 53.70 15.11 3.22 22.83 0.96 4.18 92.14 70.20 69.15 39.06 47.47 48.66 38.55 45.78 43.86
AutoScraper 62.06 14.15 3.86 15.11 0.96 3.86 91.76 78.10 76.97 56.23 27.27 67.56 53.01 34.94 60.10
[…]
Reflexion 67.50 13.75 4.37 10.94 0.94 2.50 93.28 82.76 82.40 64.81 19.51 75.85 50.60 33.73 63.50
AutoScraper 71.56 14.06 5.31 4.06 0.63 4.37 92.49 89.13 88.69 64.11 15.33 76.21 57.83 16.87 75.52
Open-source LLMs
Phi-3-medium COT 12.50 2.81 3.12 80.00 0.00 1.56 94.38 18.10 17.21 11.78 79.46 16.28 9.64 85.54 12.28
Reflexion 12.19 6.56 1.87 77.81 0.00 1.56 92.45 18.21 17.31 12.66 82.28 15.42 7.23 90.36 8.89
AutoScraper 24.06 12.50 7.50 52.19 0.31 3.44 85.07 38.59 34.93 21.15 64.42 30.29 22.89 69.88 26.60
[…]
Reflexion 18.08 4.80 2.95 73.06 0.00 1.11 78.96 23.26 22.44 13.73 80.26 16.01 8.82 85.29 12.69
AutoScraper 23.99 8.12 1.48 64.94 0.00 1.48 78.59 28.70 28.41 11.16 85.84 12.52 13.51 81.08 17.39
[…]
Reflexion 36.25 6.88 3.12 51.25 0.00 2.50 89.35 44.57 43.60 29.62 62.02 33.64 22.78 69.62 28.20
AutoScraper 46.88 10.62 7.19 30.31 0.63 4.37 87.32 62.71 59.75 40.77 38.33 52.50 36.71 43.04 48.23
[…]
Reflexion 37.19 11.25 4.06 44.69 1.25 1.56 86.41 48.28 47.08 36.24 51.92 43.64 22.89 65.06 32.04
AutoScraper 38.75 11.25 5.31 39.69 0.63 4.37 84.91 52.11 49.68 37.63 50.52 44.33 39.76 42.17 50.28
Table 2: The executable evaluation and IE evaluation of LLMs with three frameworks in Swde, Extended Swde, and Ds1 dataset. Best Correct, Unexecutable, precision, recall, and F1 score are marked bold.
4.2 Progressive Generation
Dealing with the lengthy content and hierarchical structure of webpages, generating a complete and executable scraper in one turn is difficult. However, the HTML content is organized in a DOM tree structure, which makes it possible to prune irrelevant page components and hence, limit the length and height of the DOM tree to improve the performance of LLM generation.
4.3 Synthesis
Specifically, we randomly select $n_{s}$ webpages from the case as seed webpages. Then, we generate an action sequence for each of them. Subsequently, we execute multiple different action sequences to extract information from the seed web pages, respectively. We collect all action sequences and their corresponding results and then choose one that can extract all the target information in the web pages as the final action sequence.
5 Experiment
Intending to put AutoScraper to practical use, we investigate the following research questions: 1) Can AutoScraper outperform the state-of-the-art scraper generation methods? 2) How does AutoScraper framework improve the performance of the scraper generation task? 3) Does AutoScraper meet the requirements for web scraping tasks, specifically being accurate and efficient?
5.1 Experimental Settings & Evaluation Metrics
We conduct our experiment on 8 LLMs including closed-source LLMs: GPT-3.5-Turbo OpenAI (2022), Gemini Pro Team et al. (2023), GPT-4-o-mini OpenAI (2024) and GPT-4-Turbo OpenAI (2023) as well as open-source LLMs: Phi-3-medium Abdin et al. (2024), CodeLlama-34B Rozière et al. (2024), Mixtral 8$\times$7B Jiang et al. (2024) and Deepseek-Coder-33B Guo et al. (2024). Furthermore, we apply different LLM-prompt-based web agents as our baselines, including COT Wei et al. (2023) and Reflexion Shinn et al. (2023) and AutoScraper to them. The comparison between them is discussed in Appendix B.1. Due to the limited-length context of LLMs, all experiments are conducted under zero-shot settings.
We test them on three datasets: Swde Hao et al. (2011), Extend Swde Lockard et al. (2019) and Ds1 Omari et al. (2017). The detailed experimental results of the last two can be found in Appendix A.1 and  A.2. We set the size of seed webpages $n_{s}=3$ for Swde and Extend Swde, $n_{s}=1$ for Ds1 and max retry times $d_{max}=5$.
In addition to the execution evaluation metrics described in Section 3.3, we also employ traditional evaluation metrics to more comprehensively assess the quality of different action sequences. Specifically, we adopt precision (P.), recall (R.), and macro-f1 (F1), which are calculated as the mean of the corresponding metrics for each case. Detailed experimental results on the last two datasets can be found in Table 16 and 17.
5.2 Main Results
Results in Table 4.1 show that: 1) With AutoScraper generating action sequence, LLMs can achieve better performance. Compared to the COT and Reflexion baseline, our method performs a higher ratio of correct and a lower ratio of un-executable. Also, it should be noted that Mixtral 8$\times$7B + AutoScraper can outperform GPT-3.5-Turbo + Reflexion, indicating the superiority of AutoScraper in the generation of executable action sequences in the scraper generation task. 2) Models with small parameter sizes have significant difficulties in understanding and writing executable paths, so they can be considered challenging to apply in this task. On the contrary, large-scale models demonstrate a more stable ability in instruction alignment, web structure comprehension, and reflection on execution results. 3) Traditional IE evaluation metrics cannot well describe the success rate of our task. Especially for the precision metric, it fails to reveal the performance gap among different methods with different models. This is because the extraction metrics only evaluate the results that have been extracted, ignoring that unexecutable or empty extractions also greatly damage the executability.
Models Method Exec Eval IE Eval
Correct($\uparrow$) Unex.($\downarrow$) F1
GPT-3.5-Turbo COT 36.75 43.46 47.99
   - synthesis 27.56 57.24 34.44
[…]
   - synthesis 28.62 59.01 35.01
[…]
   - synthesis 44.52 29.33 58.44
Gemini Pro COT 29.69 47.19 41.81
   - synthesis 27.56 57.24 33.09
[…]
   - synthesis 28.62 59.01 37.60
[…]
   - synthesis 39.46 31.56 56.48
GPT-4-Turbo COT 61.88 14.37 76.95
   - synthesis 46.88 30.00 61.20
[…]
   - synthesis 56.87 25.31 69.78
[…]
   - synthesis 65.31 11.87 80.41
Table 3: Ablation study on AutoScraper. We report Correct, Unexecutable from the executive evaluation, and F1 score from the IE evaluation in Swde dataset.
5.3 Ablation Study
To further justify the effectiveness of each component of AutoScraper, we perform an ablation study. The results are shown in Table 5.2. It shows that: 1) AutoScraperwithout a second module still beat the other two baseline methods among different LLMs. 2) The second module of AutoScraper, synthesis module, not only improves AutoScraper, but also improves the performance of other methods. Using more web pages for inference can make the generated scraper more stable and have better generalization.
Figure 3: The performance of AutoScraper with different number of seed websites in Swde dataset.
5.4 Seed Websites
In all previous experiments, we fixed the number of seed websites $n_{s}=3$, which demonstrates the effectiveness of the synthesis module. In this experiment, we offer different numbers of seed webpages and test the performance of AutoScraper. The result is shown in Figure 3.
As the number of seed webpages increases, the correct ratio increases, while the unexecutable ratio decreases. It suggests that the performance of AutoScraper can still be further improved by providing more seed webpages. In addition, the performance improvement reduces as the number increases, which shows that there is an upper limit to improve the performance of AutoScraper by increasing the number of seed webpages.
6 Discussion
In this section, we will discuss other aspects of AutoScraper, including its comparison with existing website information extraction methods, efficiency analysis of AutoScraper, and the limitations of the current approach.
Model Direct Extraction AutoScraper
GPT-3.5-Turbo 75.76 69.20
Gemini Pro 76.62 54.91
GPT-4-o-mini 79.93 76.97
GPT-4-Turbo 78.56 88.69
Phi-3-medium 71.73 34.93
Codellama 47.38 28.41
Mixtral 8$\times$7B 73.45 59.75
Deepseek-coder 61.96 49.68
Table 4: Comparing LLM direct extraction with AutoScraper on the Swde dataset.
6.1 Comparison with LLM Direct Extraction
Since LLMs can understand human instructions and webpage text, a natural web information extraction solution involves using prompts to guide LLMs to extract target content, which we refer to as direct extraction. We compare direct extraction with AutoScraper both in zero-shot settings using each of the LLMs mentioned above.
Table 4 shows that in the direct extraction setting, the extraction performance of all LLMs other than GPT-4-Turbo is superior to that of AutoScraper. However, as the capability of LLMs improves, the gap between the two settings narrows. This indicates that: 1. While LLMs like Phi-3-medium can understand webpage content well (i.e., correctly extract the expected content), they still struggle to comprehend webpage structures (i.e., generating XPath using features like DOM tree). 2. AutoScraper, combined with the best current LLMs, already achieves superior extraction performance, and the framework is expected to deliver even better and more stable performance as LLMs continue to improve.
Model F1
Render-Full Hao et al. (2011) 84.30
FreeDOM Lin et al. (2020) 82.32
SimpDOM Zhou et al. (2021) 83.06
[…]
WebFormer Wang et al. (2022) 86.58
Reflexion + GPT-4-Turbo 82.40
AutoScraper + GPT-4-Turbo 88.69
Table 5: Comparing the extraction performance (F1) of 5 baseline models to our method AutoScraper using GPT-4-Turbo on the Swde dataset. Each value of the supervised model in the table is trained on 1 seed site.
6.2 Comparison with supervised baselines
To further demonstrate that AutoScraper is adaptive to different web information extraction tasks, we conduct a comparison with 5 baseline models in web information extraction on supervised learning scenarios: Render-Full Hao et al. (2011) proposes a complicated heuristic algorithm for computing visual distances between predicted value nodes and adjusting the predictions. FreeDOM Lin et al. (2020) and SimpDOM Zhou et al. (2021) encode textual features of DOM tree node with LSTM, while MarkupLM Li et al. (2022) is pre-trained on HTML with text and markup information jointly. WebFormer Wang et al. (2022) leverages the web layout for effective attention weight computation. These models are trained on webpages in some seed websites and tested on the other websites.
Table 5 shows the result. Although the comparison is unfair because our method is in zero-shot settings, AutoScraper beat all of them on F1 scores. It shows that by designing an appropriate framework, LLMs can surpass supervised learning methods in some web information extraction tasks.
6.3 Efficiency Analysis
Suppose the number of seed webpages is $n_{s}$, the number of webpages on the same website is $N_{\mathcal{W}}$, the time to generate a wrapper is $T_{g}$, the time of synthesis is $T_{s}$, and the time for extracting information from a webpage with a wrapper is $T_{e}$. The total time for extracting all information from all websites with AutoScraper is
$T_{1}=T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}$ (3)
$T_{2}=N_{\mathcal{W}}T_{d}$ (4)
In a real-world scenario, there are many web pages from the same websites to be extracted. Although generating a wrapper takes more time than extracting directly from a single webpage, the extraction efficiency of subsequent web pages would be significantly improved. To explore how many webpages are needed to make AutoScraper more efficient in web IE, we calculate the threshold of $N_{\mathcal{W}}$. Suppose $T_{1}\leq T_{2}$, we have
$T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}\leq N_{\mathcal{W}}T_{d}$ (5)
$N_{\mathcal{W}}\geq\frac{n_{s}T_{g}+T_{s}}{T_{d}-T_{e}}$ (6)
Websites $T_{d}$ $n_{s}T_{g}+T_{s}$ $T_{e}$ $N_{\mathcal{W}}$
Auto 8.27s 238.4s 0.30s 30
Book 10.20s 176.4s 0.51s 18
Camera 6.59s 107.1s 0.31s 18
Job 7.42s 123.5s 0.21s 18
Movie 7.47s 133.2s 0.21s 19
Nbaplayer 8.32s 179.4s 0.45s 23
Restaurant 8.87s 160.8s 0.54s 20
University 14.26s 134.7s 0.32s 10
Table 6: Time efficiency analysis on GPT-4-Turbo.
To verify the efficiency advantages of AutoScraper in large-scale web information extraction scenarios, we conducted tests on the Swde dataset. Specifically, we randomly selected a website in each of the 10 domains. We repeat 3 times on AutoScraper and record the average time to estimate $n_{s}T_{g}+T_{s}$ and $T_{e}$. At the same time, we record the average time $T_{d}$ on 10 web pages with LLM extracting directly. We calculate the threshold of $N_{\mathcal{W}}$ following the Equation 6 and show them in Table 6. It can be observed that the threshold of the page numbers is 19.5 on average, which is significantly lower than the average number of web pages per site in Swde dataset.
6.4 Error Analysis
We perform an analysis by looking at the recorded action sequence of AutoScraper with GPT-4-Turbo and identify the following common failure modes. We mainly focus on the cases categorized as unexecutable, over-estimate, and else.
Non-generalizability of webpages
The target information and corresponding webpage structures exhibit variations across different webpages, leading to a lack of generalizability in AutoScraper (i.e., the inability to apply the same rules across all webpages in the same website). For instance, for the task "Please extract the name of the company offering the job" in the website job-careerbuilder, most webpages contain the company name, but there is one webpage where the company name is "Not Available" on another node of DOM tree.
Miss in multi-valued
Presented with the task of generating a scraper for extracting address in restaurant webpages or contact phone number from university websites, the target information is located in multiple locations in the webpage, such as the information bar, title, etc. Although AutoScraper is capable of generating action sequences to extract portions of information, crafting a comprehensive action sequence that captures all of the information remains a challenge.
7 Conclusion
In this paper, we introduce the scraper generation task and the paradigm that combines LLMs and scrapers to improve the reusability of the current language-agent-based framework. We then propose AutoScraper , a two-phase framework including progressive generation and synthesis module to generate a more stable and executable action sequence. Our comprehensive experiments demonstrate that AutoScraper can outperform the state-of-the-art baseline in the scraper generation task.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 62102095). The computations in this research were performed using the CFFF platform of Fudan University. The authors would like to express their sincere gratitude to Alibaba (China) Co., Ltd. and Alibaba Holding-Aicheng Technology-Enterprise Intelligence Business Unit for their support.
Limitation
We introduce a paradigm that combines LLMs with scrapers for web scraper generation tasks and propose AutoScraper to generate an executable action sequence with progressively understanding the HTML documents. Though experimental results show the effectiveness of our framework, there are still some limits to our work.
First, our framework is restricted to the paradigm in the information extraction task for vertical web pages. LLMs with scrapers provide high efficiency in open-world web IE tasks, but can hardly transfer to existing web environments such as Mind2Web Deng et al. (2023), WebArena Zhou et al. (2023).
Second, our framework relies on the performance of backbone LLMs. Enhancing LLMs’ ability to understand HTML is a very valuable research question, including corpus collection and training strategy. We will research HTML understanding enhancement in future work.
Ethic statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honour the code of conduct.
Use of Human Annotations
Risks
Appendix A Experiments
Models Method Executable Evaluation IE Evaluation
Closed-source LLMs
GPT-3.5-Turbo COT 35.19 3.48 4.53 55.40 0.35 1.05 88.66 42.86 41.28
Reflexion 43.90 1.74 2.09 49.13 0.35 2.79 93.46 49.58 48.66
AutoScraper 46.34 4.18 8.01 34.84 0.35 6.27 84.65 61.88 57.74
[…]
Reflexion 34.15 2.09 6.97 51.57 0.35 4.88 84.43 45.19 41.66
AutoScraper 35.89 5.23 10.10 42.86 0.35 5.57 83.80 52.83 47.80
[…]
Reflexion 39.06 7.07 2.02 47.47 0.34 4.04 95.03 49.07 48.66
AutoScraper 56.23 5.39 5.05 27.27 0.00 6.06 91.12 69.45 67.56
[…]
Reflexion 64.81 4.18 5.57 19.51 0.35 5.57 87.39 77.81 75.85
AutoScraper 64.11 3.48 6.27 15.33 0.35 10.45 82.71 80.25 76.21
Open-source LLMs
Phi-3-medium COT 11.78 1.01 5.05 79.46 0.34 2.36 91.03 19.08 16.28
Reflexion 12.66 1.90 1.90 82.28 0.00 1.27 93.87 16.03 15.42
AutoScraper 21.15 2.88 7.69 64.42 0.00 3.85 87.88 33.39 30.29
[…]
Reflexion 13.73 1.72 3.00 80.26 0.00 1.29 89.41 17.76 16.01
AutoScraper 11.16 0.00 1.72 85.84 0.00 1.29 92.49 13.29 12.52
[…]
Reflexion 29.62 1.05 4.18 62.02 0.35 2.79 83.44 36.44 33.64
AutoScraper 40.77 3.83 9.76 38.33 0.35 6.97 82.50 58.14 52.50
[…]
Reflexion 36.24 3.48 3.83 51.92 0.00 4.53 83.53 45.03 43.64
AutoScraper 37.63 2.44 5.92 50.52 0.35 3.14 86.91 47.09 44.33
Models Method Executable Evaluation IE Evaluation
Closed-source LLMs
GPT-3.5-Turbo COT 32.65 4.08 8.16 53.06 0.00 2.04 90.56 43.54 41.16
Reflexion 36.73 8.16 4.08 51.02 0.00 0.00 95.56 44.22 43.75
AutoScraper 48.98 4.08 0.00 44.90 0.00 2.04 94.90 51.70 52.38
[…]
Reflexion 20.25 10.13 1.27 65.82 0.00 2.53 88.83 26.93 27.66
AutoScraper 43.04 15.19 3.80 34.18 0.00 3.80 93.76 55.97 56.92
[…]
Reflexion 38.55 13.25 2.41 45.78 0.00 0.00 91.40 43.68 43.86
AutoScraper 53.01 6.02 4.82 34.94 0.00 1.20 79.06 61.03 60.10
[…]
Reflexion 50.60 10.84 4.82 33.73 0.00 0.00 96.85 62.65 63.50
AutoScraper 57.83 15.66 4.82 16.87 0.00 4.82 92.88 74.95 75.52
Open-source LLMs
Phi-3-medium COT 9.64 4.82 0.00 85.54 0.00 0.00 95.18 11.76 12.28
Reflexion 7.23 0.00 1.20 90.36 0.00 1.20 97.87 9.47 8.89
AutoScraper 22.89 3.61 3.61 69.88 0.00 0.00 88.00 28.22 26.60
[…]
Reflexion 8.82 0.00 5.88 85.29 0.00 0.00 94.12 14.41 12.69
AutoScraper 13.51 0.00 5.41 81.08 0.00 0.00 84.12 18.92 17.39
[…]
Reflexion 22.78 6.33 1.27 69.62 0.00 0.00 94.15 28.03 28.20
AutoScraper 36.71 11.39 6.33 43.04 0.00 2.53 91.59 48.52 48.23
[…]
Reflexion 22.89 6.02 3.61 65.06 0.00 2.41 90.21 31.43 32.04
AutoScraper 39.76 10.84 6.02 42.17 0.00 1.20 90.43 51.39 50.28
Table 8: The executable evaluation and IE evaluation of LLMs with three frameworks in Ds1 dataset. We examine 8 LLMs, including 4 closed-source LLMs and 4 open-source LLMs.
A.1 Main results on Extended Swde
Table A shows the result. By comparing Table 4.1, we find that: 1) Under complex extraction task settings (multiple target values and ambiguous problem description), the closed-source LLMs perform better in generating executable action sequences compared to the open-source LLMs. 2) There are some tasks with unclear descriptions, such as the "Calendar System" and "Facilities and Programs Offered" on university websites, which affect the wrapper generation performance of all methods.
A.2 Main results on Ds1
Due to Ds1 only contains 166 hand-crafted webpages, and for each website, there are only two webpages, so we take one webpage for inference and the other for evaluation. Meanwhile, due to the number of seed websites equal to one, we test three methods without applying the synthesis module described in Section 4.3.
Table A shows the result in the DS1 dataset. Among all LLMs with three methods, GPT-4-Turbo + AutoScraper achieves the best performance, and AutoScraper beats the other two methods in all LLMs, which is consistent with our conclusion.
A.3 Generate with Golden Label
Models Method Executable Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else
Closed-source LLMs
GPT-3.5-Turbo COT 41.70 12.92 7.38 35.42 0.74 1.85
Reflexion 47.23 16.24 2.21 33.21 0.37 0.74
AutoScraper 56.89 19.43 5.65 13.43 0.71 3.89
[…]
Reflexion 35.31 9.38 6.88 43.75 1.56 3.12
AutoScraper 45.31 13.44 6.25 30.31 1.25 3.44
[…]
Reflexion 62.38 10.29 1.61 23.15 0.64 1.93
AutoScraper 67.20 12.22 3.86 12.22 0.96 3.54
[…]
Reflexion 71.25 7.19 4.69 14.37 0.94 1.56
AutoScraper 75.31 10.94 4.37 4.06 0.63 4.69
Open-source LLMs
Phi-3-medium COT 11.11 4.13 1.27 82.22 0.00 1.27
Reflexion 12.19 5.27 7.59 72.43 0.31 2.21
AutoScraper 27.27 16.45 9.52 41.56 0.87 4.33
[…]
Reflexion 22.21 4.93 3.94 66.95 0.49 1.48
AutoScraper 26.20 12.55 5.54 53.51 0.00 2.21
[…]
Reflexion 34.69 8.13 5.31 49.06 0.63 2.19
AutoScraper 45.62 11.56 5.94 32.50 1.25 3.12
[…]
Reflexion 38.75 11.87 2.81 42.19 0.63 3.75
AutoScraper 38.44 20.94 4.06 31.56 0.94 6.56
Table 9: The executable and IE evaluation with 8 LLMs on Swde dataset with golden label.
To better illustrate the effectiveness of our framework in generating executable action sequences, we compare the performance of COT, Reflexion, and AutoScraper , while answering the instruction. By offering the same extraction targets, we can effectively detect the performance of different frameworks in generating action sequences.
Table A.3 shows experimental results, from which we can have the following observations: 1) Our proposed progressive understanding framework still effectively enhances the model’s performance under this setting; 2) LLMs still suffer in accurately understanding web page contents with semi-structured markup languages, which illustrate the performance gap between Table 4.1 and Table A.3; 3) Compared to closed-source LLMs, even provided with golden labels, Open-source LLMs are unable to achieve sustained performance improvement. This phenomenon demonstrates that the bottleneck for these models lies not in understanding the webpage content but in understanding the webpage’s hierarchical structure itself.
@@ 18       $k\leftarrow k+1$;
Appendix B Analysis on AutoScraper
B.1 Comparison with COT & Reflexion
Figure 4 more intuitively shows the specific differences between different baselines in the experiment. The most significant difference between AutoScraper and other methods lies in whether the hierarchical structure of web pages is utilized to help LLMs reduce the difficulty of complex web structures. COT only executes one turn while the other executes multiple turns and can learn from the failed execution of the wrapper. Compared to the Reflexion method, AutoScraper employs top-down and step-back operations to prune the DOM tree during each XPath generation process, thereby reducing the length of the web page. In contrast, the Reflexion method can only reflect and regenerate after producing an unexecutable XPath, which does not effectively simplify the webpage.
Figure 4: Comparison of AutoScraper with COT and Reflexion.